OpenStudy (anonymous):

Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem. Can someone show me how to set this problem up? I'm really confused on what to do. If someone would just set up the problem for me I'd REALLY appreciate it!

OpenStudy (anonymous):

i fan and medal too!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

they sell a flower at 140% from the original price it would be 0.50$......100% x$.............140% x= 0.50*140/100=0.70$ a flower to make 500$ you should sell 500/0.70= 714 flowers

OpenStudy (anonymous):

Thanks for answering! But i read a bunch of answers for the same problem on open study and they all came up with 2500 or something. Are you sure its 714?

OpenStudy (anonymous):

i think i made a mistake

OpenStudy (anonymous):

OH ok i get it now

OpenStudy (anonymous):

Okay can you show me like step by step because im really slow lol

OpenStudy (anonymous):

P = the profit they make n = the number of flowers they sell m = the mark-up of the flowers c = the cost of a flower (PNMC)

OpenStudy (anonymous):

then P = mcn - cn

OpenStudy (anonymous):

Got it

OpenStudy (anonymous):

Then what

OpenStudy (anonymous):

rearrange to find n P = n(mc - c) n = P / (mc - c)

OpenStudy (anonymous):

substituting values n = 500 / (1.4*0.5 - 0.5) n = 2500 they must be selling 2500 flowers to make a profit of $500

OpenStudy (anonymous):

Ohh, thank you so much this really helped!!!!!!!

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

:)